Statemere ษd Report Form of Balance Sheet he tolowing selected accounts and their cument buliances a ppear in the ledger of Kanour Co. for the fiscal year ended une 30,2016 92,000 Divideds 50,000 Sales 375000 Cost of Merchandise Sold 0,000 Sales Salaries Expense Prepaid Insuranc Office Equimen 2000 dvertising Expense 3,000 18,000 tone Equipment 50,000 Office Saleres Expense 7,500 Rent Expee 8,500 nsuance Expense 4,000 40000 Office Supplies Expes Salaries Payable oe Payable(napmu 2032) 10,000 81,000 nterest Expense 5 6
Answer #1
A B C D E F G
2
3 Stepwise income statement can be prepared as follows:
4
5 Net Sales $8,925,000
6 Cost of goods Sold $5,620,000
7 Gross margin $3,305,000 =E5-E6
8 Operating expenses:
9 Sales Salaries Expenses $850,000
10 Advertising Expense $420,000
11 Depreciation Expense - store equipment $33,000
12 Miscellaneous Selling Expense $18,000
13 Office Salaries Expense $540,000
14 Rent Expense $48,000
15 Insurance Expense $24,000
16 Depreciation Expense - Office Equipment $10,000
17 Office Supplies Expense $4,000
18 Miscellaneous Administrative Expense $6,000
19 Operating Expense $1,953,000 =SUM(D9:D18)
20 Operating income (EBIT) $1,352,000 =E7-E19
21 Other income (Expense)
22 Interest Expense ($12,000)
23 Total Other Income ($12,000)
24 Taxable income $1,340,000
25 Tax Expense $0
26 Net income $1,340,000 =E24-E25
27
28
29
30 Report form of Balance sheet will be as follows:
31
32 June 30 2016
33 Assets
34 Cash $92,000
35 Account Receivables $450,000
36 Merchandise Inventory $375,000
37 Office Supplies $10,000
38 Prepaid Insurance $12,000
39 Current Assets $939,000 =SUM(D34:D38)
40 Office Equipment $220,000
41 Less: Accumulated Depreciation - Office Equipment ($58,000)
42 Store Equipment $650,000
43 Less: Accumulated Depreciation - Store Equipment ($87,500)
44 Net Fixed Assets $724,500 =SUM(D40:D43)
45 Total Assets $1,663,500 =D39+D44
46
47 Liabilities and shareholders' equity
48 Account Payable $48,500
49 Salaries Payable $4,000
50 Current Liabilities $52,500 =SUM(D48:D49)
51 Notes Payable (Long Term) $140,000
52 Total Liabilities $192,500 =D50+D51
53 Common Stock $50,000
54 Retained Earnings $1,721,000 =381000+E26
55 Dividends ($300,000)
56 Shareholders' equity $1,471,000 =SUM(D53:D55)
57 Total Liabilities and shareholders' equity $1,663,500 =D52+D56
58